OpenStudy (anonymous):

PLLEEAASEE NEED HELP NOW!!! Graph x + 3y ≤ 6, indicating the solution set with crosshatching or shading. Explain how you determined where to draw the line and shade the area that represents the solution set.

OpenStudy (daniellelovee):

you have to make the equation equal y=mx+b using PEMDAS backwards

OpenStudy (anonymous):

I don't understand.. can you walk me through it?

OpenStudy (daniellelovee):

sure (I just can't give you the answer because I have a warning)

OpenStudy (daniellelovee):

ok so you have to imagine that the inequality symbol is an equal sign therefore if you had to solve for y in the equation of x+3y=6 how would you solve it?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you would first divide everything out by 3 and you would get x/2+y=2

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (daniellelovee):

no remember that you are using PEMDAS backwards so you would substract x first

OpenStudy (daniellelovee):

P-parenthesis E-exponent M-multiplication D-division A-addition S-substraction

OpenStudy (anonymous):

oh ok so 3y=6-x

OpenStudy (daniellelovee):

yes and now you are just missing one more step to get your equation

OpenStudy (anonymous):

now divide by 3 so it would be y=2-x/3

OpenStudy (daniellelovee):

YES! :) you got it and now since we are using slope standard form we would change y=2-x/3 to y=-3/x+2

OpenStudy (daniellelovee):

and put the inequality symbol into the = sign

OpenStudy (anonymous):

how would I graph it?

OpenStudy (daniellelovee):

have you graphed y=mx+b before? you know x is the slope and +2 is the y intercept

OpenStudy (anonymous):

yes

OpenStudy (daniellelovee):

ok then just do that until you have a line and then I'll tell you how to shade :)

OpenStudy (anonymous):

Thanks so much you're a life saver :)))))

OpenStudy (daniellelovee):

haha np ^.^

OpenStudy (daniellelovee):

so you have your graph? :)

OpenStudy (daniellelovee):

btw just a tip, since the slope is negative your line will go down